OpenStudy (anonymous):

What is the volume of a cone with a radius of 2 and a height of 19? A. 79.59 units3 B. 238.76 units3 C. 39.79 units3 D. 722 units3

OpenStudy (anonymous):

@jakashaka123

OpenStudy (anonymous):

volume of the cone is= (pi x r^2)H/3 therefore if r=2 H=19 pi=3.14 so the answer is A=79.59 units3

OpenStudy (anonymous):

The formula for volume of cone is V = ( h x A)/3 , where h is the height of the cone A is the area of the base , which is equal to A = pi x r^2 , where r is the radius V = (h x pi x r^2) / 3 Assume pi = 3.1416 V = (19 x 3.1416 x (2)^2 )/3 V = (19 x 3.1416 x 4) /3 V = 238.762/3 V = 79.59 units 3

OpenStudy (anonymous):

\[v=1/3 \pi r^3h\] \[1/3 \times \pi \times 2^3 \times 19\] can u solve it further

OpenStudy (anonymous):

...

OpenStudy (anonymous):

u just had to do that didnt you